Grade 9General Physics

A body travels 2m in the 2nd second and 6m in the next 4 seconds.What will be the distance travelled in the 9th second?

Sanskar sahoo
9 Years agoGrade 9

1 Answer

RiQ
ApprovedApproved Tutor Answer9 Years ago
S= u + ½a(2n-1).S=distance covered in that second.U=initial speed.a= acceleration.n= second .2=u + ½a(4-1) ........(i),6=u + ½a(8-1) ........(ii).U=-1m/s. a=2m/s². S= -1 +½{2(18-1)} = 16m . ✓For more help visit RiQ UR Classes on YouTube. Thanks!